A solution containing 2.675 g of 33 6NHCoCl · (molar mass = 267.5 g 1 mol- ) is passed through a cation
exchanger . The chloride ions obtained in solution were treated with excess of 3AgNO to give 4.78 g of
AgCl (molar mass =143.5 g 1 mol- ). The formula of the complex is

Grade:12

1 Answers

Gaurav
9 years ago
Energy of an electron in the ground state = -2.18 x 10-18.
Energy required for removal of the electron = 0 -(-2.18 x 10-18 ) =2.18 x 10-18 J.
This is the energy required to remove the electron from one atom of hydrogen in ground state. The energy required to remove electron from one mole (6.022 x 1023atoms) will be = 2.18 x 10-18 x 6.022 x 1023 = 13.12 x 105 J/mol